OpenStudy (anonymous):

What is the sum of the multiples of 4 from 16 to 100?

OpenStudy (anonymous):

\[\sum_{4}^{20}4x=\frac{ 4(16(16+1)) }{ 2 }=544\]

OpenStudy (anonymous):

the answer choices are 1,218 1,276 1,300 1,334

OpenStudy (rosedewittbukater):

I'm doing this test right now. I have no idea if I'm right because I havent finished it yet, but I did 25-3 to get 22 because thats how many numbers are in between 4 and 25 including 4 and 25. Then I used the formula the book gave us \[S _{n}=\frac{ n }{ 2 }(a _{1}+a _{n)}\] \[S _{22}=\frac{ 22 }{ 2 }(116)\] \[=\frac{ 2552 }{ 2 }\] \[=1276\] At first I just did it in a calculator by doing (4x4)+(4x5)+(4x6)+(4x7) all the way to 4x25 and got the same thing. Then I tried the equation. Let me know if you submitted it and what you got. This unit is confusing me a lot so far.